OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8>

OpenStudy (anonymous):

They're parallel. The "gradient"=y/x=-2 for both vectors.

OpenStudy (ipwnbunnies):

A test for orthogonal is to take the dot product of the two vectors. If the dot product equals 0, then the two vectors are orthogonal.

OpenStudy (anonymous):

\[u=1i-2j,v=-4i+8j\] \[u.v=\left( i-2j \right).\left( -4i+8j \right)=1\left( -4 \right)+\left( -2 \right)\left( 8 \right)\] =-4-16 \[=-20\neq 0\] hence they are not orthogonal.

OpenStudy (ipwnbunnies):

Just saying lol.

OpenStudy (anonymous):

thnk u lool

OpenStudy (anonymous):

i hve another question Write the complex number in the form a + bi. 6(cos 330° + i sin 330°)

OpenStudy (anonymous):

For a unit circle: 330 degrees is basically 330-360=-30 degrees. Now I think you can do the rest. :D